OpenStudy (anonymous):

Which equation describes this statement? The sum of c and d is equal to 28. Let d = 7

Mehek (mehek14):

where are the equations?

OpenStudy (anonymous):

c ÷ 7 = 28 B. c × 7 = 28 C. c – 7 = 28 D.c+7=28

Mehek (mehek14):

so which one do you think it is?

Mehek (mehek14):

do you know what sum means?

OpenStudy (jjuden):

but it say sum sum means +

OpenStudy (jjuden):

so your answer is D

OpenStudy (anonymous):

i feel so stupid XD

Mehek (mehek14):

@jjuden this isn't your question so don't answr when im asking t

Mehek (mehek14):

@creepypastaXD do you know what sum means?

OpenStudy (anonymous):

yea

OpenStudy (jjuden):

im not answering what your asking im helping them so dont get all b*tchy

Mehek (mehek14):

@creepypastaXD so what sign would it bfor sum? + - / or *

OpenStudy (anonymous):

* i guess

Mehek (mehek14):

no not multiplication then it would be product

OpenStudy (jjuden):

no sum means +

OpenStudy (anonymous):

p.s what do you think of my name

Mehek (mehek14):

that is irrelevant to the question @creepypastaXD

OpenStudy (jjuden):

The sum of c and d is equal to 28. Let d = 7 so c+7=28

OpenStudy (jjuden):

@creepypastaXD love the name :)

OpenStudy (anonymous):

thanks im like in love with creepypasta

OpenStudy (anonymous):

do you know what it is?

OpenStudy (jjuden):

no

OpenStudy (anonymous):

warning : crazy!!!! pyco killers

OpenStudy (anonymous):

you think im insane don't you??